The probability that a given tourist goes to the amusement park is 0.47, and the probability that she goes to the water park is 0.58. If the probability that she goes to either the water park or the amusement park is 0.75. What is the probability that she visits both of the parks on vacation? 2.
Answer #1

P(amusement park) = 0.47 , P(water park) = 0.58

P( amusement park OR water park) = 0.75

P( amusement park AND water park) = ?

Using addition rule of probability,

P(A OR B) = P(A) + P(B) - P(A AND B)

0.75 = 0.47 + 0.58 - P( amusement park AND water park)

Therefore,

P( amusement park AND water park) = 0.47 + 0.58 - 0.75

P( amusement park AND water park) = 0.30

Probability that she visits both of the parks on vacation = 0.30